# Decoding Intelligence: A Framework for Certifying Knowledge Comprehension in LLMs

### v2

@@ Abstract
Knowledge comprehension capability is an important aspect of human intelligence. As Large Language Models (LLMs) are being envisioned as superhuman agents, it is crucial for them to be proficient at knowledge comprehension. However, existing benchmarking studies do not provide consistent, generalizable, and formal guarantees on the knowledge comprehension capabilities of LLMs. In this work, we propose the first framework to certify knowledge comprehension in LLMs with formal probabilistic guarantees. Our certificates are quantitative — they consist of high-confidence, tight bounds on the probability that a target LLM gives the correct answer on any knowledge comprehension prompt sampled from a distribution. We design and certify novel specifications that precisely represent distributions of knowledge comprehension prompts leveraging knowledge graphs. We certify SOTA LLMs for specifications over the Wikidata5m knowledge graph. We find that the knowledge comprehension capability improves significantly with scaling the size of the models.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated human-level performance for several real-world tasks (Street et al., 2024; Yang et al., 2023b; Bommasani et al., 2022; Harrison, 2024). An important use case for LLMs is knowledge comprehension (Lazaridou et al., 2022; Khattab et al., 2023), that is, they are often used to summarize long texts and webpages (Perplexity, 2023), respond to user queries based on the context (Yang et al., 2018), and serve as adaptive task decomposers for reasoning-based retrieval augmented generation tasks (Yao et al., 2023b). Knowledge comprehension consists of answering questions by extracting relevant information from large, unstructured texts and reasoning with it. Large context windows of millions of tokens in models like Gemini v1.5 (Gemini Team, 2024) reduce reliance on large knowledge corpora of RAG systems and parametric knowledge held by LLMs. Users increasingly provide extensive references during inference to guide responses. This makes analyzing the knowledge-comprehension and reasoning capabilities of popular LLMs crucial. Moreover, knowledge comprehension is considered a basic evaluation of language understanding in human learners, according to the Bloom’s taxonomy (Bloom, 1956). Students are tested on knowledge comprehension tasks at all levels of school education (National Center for Education Statistics, 2024). Standardized tests such as TOEFL (Educational Testing Service, 2024) and IELTS (IDP IELTS, 2024) contain knowledge comprehension as entire assessments. As LLMs are envisioned to become superhuman agents (Xi et al., 2023), it is imperative to formally analyze them on tasks on which humans are extensively tested, like knowledge comprehension.
There are several benchmarking studies on the performance of LLMs for knowledge comprehension (Liang et al., 2023; Chen et al., 2021; Yang et al., 2018; Wang et al., 2023a; Trivedi et al., 2022; Tang & Yang, 2024). Several of these studies use multi-hop question-answering datasets that consist of questions requiring several sequential reasoning steps to obtain the final correct answer. Thus, benchmarking knowledge comprehension often involves analyzing whether the target LLM can combine multiple pieces of information in meaningful ways and reason its way to the correct answer in the prompt, without deviating or hallucinating. However, the empirical nature of prior studies results in inconsistency in their observations (Wei et al., 2023b; Olsson et al., 2022; Shi et al., 2024). Moreover, while they can convey some high-level trends in the performance of popular LLMs, the results are devoid of any formal guarantees on their applicability. Such guarantees are crucial when deploying LLMs in large-scale knowledge-comprehension tasks in critical domains such as medicine or finance, as they give more confidence about the trustworthiness of LLMs before deployment. Our work aims to bridge this gap by introducing a novel formal certification method, QuaCer-C [^1] Quantitative Certification of Knowledge Comprehension, for obtaining certified bounds on the knowledge comprehension capability of LLMs.
Key challenges. We face the following challenges when developing a formal certification framework for knowledge comprehension in LLMs. (1) We need formal representations capturing the knowledge comprehension property, amenable to certification. Such representations should precisely capture large, diverse sets of prompts (created by varying questions, supporting texts, etc.) for knowledge comprehension and their correct responses. (2) Failure examples where the desirable property does not hold are fairly easy to construct for LLMs by appropriate prompt tuning (Xu et al., 2024; Vega et al., 2023), making binary certificates that indicate whether an LLM satisfies a specification trivially false. (3) Giving provable guarantees on LLMs is a hard, open problem, due to the high number of parameters of the models, for which the traditional certifiers (Singh et al., 2019; Shi et al., 2020; Bonaert et al., 2021) would lose significant precision leading to inconclusive analysis. The number of prompts over which we desire the target LLMs to reason correctly is also large, making enumeration-based methods for formal guarantees infeasible.
Our approach. We conceptualize the knowledge comprehension property as a novel formal specification using a knowledge graph. Instead of specifying correctness of LLM responses for all prompts in any given set, we specify a quantitative property, which is the probability of correct response for any knowledge comprehension prompt sampled from a distribution developed using a given knowledge graph. We propose a model-agnostic, black-box quantitative certification approach, QuaCer-C, which circumvents the issues that traditional approaches have with the number of parameters in LLMs and can even work for closed-source API-access LLMs. QuaCer-C generates high-confidence bounds on the quantitative property using queries, leveraging binomial proportion confidence intervals (Clopper & Pearson, 1934). Estimating with bounds is beneficial as they account for the uncertainty in the estimation as well. While formal analysis has been conducted on the individual generations of LLMs in prior work (Quach et al., 2024), there is no analysis for the average-case risk of LLMs in knowledge comprehension. Our certificates contain provable bounds on the probability of getting correct responses for any random knowledge comprehension task sampled from the distributions given in the specifications.
Contributions. We make the following contributions:
@@
    We specify the knowledge comprehension property desirable from the LLM responses as a formal specification. Our specifications use popular knowledge graphs such as Wikidata5m (Wang et al., 2021) that are augmented with supporting information about each of their entities. The specifications represent a large set of knowledge comprehension prompts with their respective correct answers expected from any target LLM.
@@
    We model certification in a target LLM as a probability estimation problem and leverage Clopper-Pearson confidence intervals to generate provable, high-confidence bounds on the quantitative property of interest. Our implementation is provided at https://anonymous.4open.science/r/QuaCer_CAnon-4130.
@@
    We generate the proposed certificates for the popular LLMs: Llama-3, Mistral, Phi-3, GPT-4o, and Gemini-1.5-Pro. We observe that as the number of model parameters increases, the knowledge comprehension capability of the LLM improves. On comparing different model classes, we see Phi-3 models performing the best among the smaller, open-source models.
Our work is the first step towards providing guarantees on the knowledge comprehension and reasoning capabilities of LLMs, to ameliorate the caution needed when using LLMs (Shanahan, 2023) in a systematic way. We anticipate it to go a long way in making LLMs faithful to traditional notions of reasoning, and hence more trustworthy for deployment in critical domains.
2 Background
2.1 Knowledge graph
A knowledge graph ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ is a collection of nodes ${\mathcal{N}}$ representing entities interconnected by directed edges in ${\mathcal{E}}$ representing their relations (Peng et al., 2023; Ji et al., 2022). They are commonly used in search engines to enhance the relevance of responses to user queries.
Figure 1: A subgraph of Wikidata5m
Hence, major companies develop their own closed-source knowledge graphs. Wikidata5m (Wang et al., 2021), a popular open-source knowledge graph consisting of 5M nodes, is a structured representation of Wikipedia pages. Each Wikidata node corresponds to a Wikipedia page, containing the page’s abstract and a set of aliases that can be synonymously used to refer to the node. Two nodes $(v_{1},v_{2})$, $v_{1},v_{2}\in{\mathcal{N}}$ are connected by a labeled edge if there is a link in the supporting document for $v_{1}$ to that for $v_{2}$. The edge $(v_{1},v_{2})$ is labeled by a set of aliases to describe the relation between the nodes. Figure 1 shows a subgraph of Wikidata5m.
2.2 Large Language Models
Large Language Models (LLMs) are autoregressive causal language models that operate on a vocabulary $\mathcal{V}$, a set of tokens. LLMs takes a sequence of tokens $x_{1},...,x_{n}$ where $x_{i}\in\mathcal{V},n>0$ and outputs a probability distribution over $\mathcal{V}$ for the potential next token $x_{n+1}$. These models are typically pretrained on vast corpora of text data  (Liu et al., 2024) and have shown remarkable capabilities  (Touvron et al., 2023; Gemini Team, 2024; OpenAI, 2024). Numerous benchmarks  (Yang et al., 2018; Rein et al., 2023; Hendrycks et al., 2021) have been developed to evaluate the performance of LLMs on tasks related to multi-step reasoning, knowledge comprehension and question answering. However, there remains a gap in our theoretical understanding of LLMs’ capabilities.
2.3 Information extraction and reasoning
Information extraction (IE) and reasoning are important research problems in natural language processing. IE involves “extracting structured information from unstructured or semi-structured data” (Chen et al., 2022) such as textual documents. Examples of IE are event extraction (Wadden et al., 2019) and relationship extraction (Pawar et al., 2017). Reasoning is the ability of a model to connect multiple facts using the correct logical operations to arrive at a final answer (Huang & Chang, 2023; Zhang et al., 2024). Typically, reasoning capabilities of LLMs are enhanced by using techniques such as Chain of Thought reasoning and its variants (Wei et al., 2023a; Yao et al., 2023a; Wang et al., 2023b), using world models (Hao et al., 2023), etc. It is evaluated in several tasks such as planning (Wang et al., 2024), mathematical reasoning (Imani et al., 2023), commonsense reasoning (Zhao et al., 2023), etc.
@@ 3 Certifying knowledge comprehension
Knowledge comprehension is the ability of a model to accurately reason through a multi-hop question (Yang et al., 2018) (combination of multiple simple information extraction questions that should be answered sequentially to arrive at the final answer) and extract the answers to intermediate questions from information provided in its prompt to reach the correct final answer. Thus, knowledge comprehension is a combination of reasoning and information extraction. Figure 2 gives an overview of our certification framework, QuaCer-C. We formally define knowledge comprehension using a knowledge graph ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ (Section 2.1) next. Our framework is agnostic to the internal structure of the target model ${\mathcal{L}}$ which can be any text-to-text generating model.
  Knowledge Graph Grammar   1. $\gamma$    := $\mathcal{V}^{+}$ 2. $\eta$    := $\mathcal{V}^{+}$ 3. $\mathcal{A}$    := $[\eta_{1},\eta_{2},\dots]$ 4. $v$    := $(\gamma,\mathcal{A})$ 5. $e$    := $((v_{1},v_{2}),\mathcal{A})$ 6. $\mathcal{N}$    := $[v_{1},v_{2},\dots]$ 7. $\mathcal{E}$    := $[e_{1},e_{2},\dots]$ 8. $\mathcal{G}$    := $(\mathcal{N},\mathcal{E})$
Figure 2: Overview of QuaCer-C. (a) A knowledge graph ${\mathcal{G}}$ pivoted on some node $v_{1}$, in this case on ‘Paracetamol Overdose’. (b) A randomly chosen path $\Pi$ originating at $v_{1}$ from the various other possible paths from $v_{1}$ in ${\mathcal{G}}$. (c) A prompt created from $\Pi$ having contexts of the nodes in $\Pi$, a distractor context (highlighted in orange, as the node for ‘Vomiting’ is a distractor for $\Pi$), and a query from $\Pi$. (d) The target LLM’s response to the prompt, validated using the correct answer. (e) Certifier obtains bounds on the probability of correct response using $n$ samples of LLM responses.
@@ 3.1 Specifying knowledge comprehension
The adjacent grammar defines a knowledge graph. Let $\mathcal{V}$ denote the vocabulary of ${\mathcal{L}}$’s tokens. $\mathcal{V}^{+}$ denotes the set of concatenation of non-empty sequences of elements of $\mathcal{V}$. Let each node $v$ of ${\mathcal{G}}$ (line 3) consist of a finite list of synonymous names (a.k.a. aliases, $\mathcal{A}$) that can be used to refer to the node, and a context $\gamma$ that provides more information about the node and its relations with other nodes. For example, in Wikidata5m (Wang et al., 2021), each node has aliases consisting of the identifiers mentioned for the subject of the corresponding Wikipedia page and context which is the abstract of the page. Each edge $e$ in ${\mathcal{G}}$ (line 3) is an ordered pair of related nodes where the relation is identified by a set of synonymous aliases $\mathcal{A}$ for the edge. Let $(v_{1},v_{2})$ denote any edge between nodes $v_{1}$ and $v_{2}$ in ${\mathcal{G}}$. We define ${\mathcal{G}}$ (line 3) as a finite collection of nodes ${\mathcal{N}}$ and edges ${\mathcal{E}}$. A path in ${\mathcal{G}}$ (Definition 3.1) is a set of connected nodes in ${\mathcal{N}}$.
@@
(Path in a Knowledge Graph). A path $\Pi=[v_{1},v_{2}\dots,v_{l}]$ is an ordered collection of nodes in a given knowledge graph ${\mathcal{G}}$, where $l>1$, such that $\forall i\in[l-1],v_{i}\in{\mathcal{N}}$, $(v_{i},v_{i+1})\in{\mathcal{E}}$, and $\forall j\in[l],i\neq j\implies v_{i}\neq v_{j}$. $\Pi_{H}\coloneqq v_{1}$ and $\Pi_{T}\coloneqq v_{l}$ are the head and tail nodes respectively of $\Pi$. Let the $i^{th}$ ($i\in[1,l]$) nodes of $\Pi$ from $\Pi_{H}$ and backwards from $\Pi_{T}$ be $\Pi[i]\coloneqq v_{i}$ and $\Pi[-i]\coloneqq v_{l-i+1}$ respectively.
Definition 3.2 describes a multi-hop reasoning problem, derived from a given knowledge graph ${\mathcal{G}}$. As ${\mathcal{G}}$ naturally encodes several multi-hop problems, we use it to form the specification for a target language model ${\mathcal{L}}$, similar to prior works such as (Ho et al., 2020; Jiang et al., 2023b).

### v3

# Certifying Knowledge Comprehension in LLMs

@@ Abstract
Large Language Models (LLMs) are increasingly deployed in safety-critical systems where they provide answers based on in-context information derived from knowledge bases. As LLMs are increasingly envisioned as superhuman agents, their proficiency in knowledge comprehension—extracting relevant information and reasoning over it to answer questions, a key facet of human intelligence—becomes crucial. However, existing evaluations of LLMs on knowledge comprehension are typically conducted on small test sets, but these datasets represent only a tiny fraction of the vast number of possible queries. Simple empirical evaluations on these limited test sets raises concerns about the reliability and generalizability of the results.
In this work, we introduce the first specification and certification framework for knowledge comprehension in LLMs, providing formal probabilistic guarantees for reliability. Instead of a fixed dataset, we design novel specifications that mathematically represent prohibitively large probability distributions of knowledge comprehension prompts with natural noise, using knowledge graphs. From these specifications, we generate quantitative certificates that offer high-confidence, tight bounds on the probability that a given LLM correctly answers any question drawn from the specification distribution. We apply our framework to certify SOTA LLMs in two domains: precision medicine and general question-answering. Our results reveal previously unrecognized vulnerabilities in SOTA LLMs due to natural noise in the prompts. Additionally, we establish performance hierarchies with formal guarantees among the SOTA LLMs, particularly in the context of precision medicine question-answering.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated human-level performance in various real-world tasks (Street et al., 2024; Yang et al., 2023; Bommasani et al., 2022; Harrison, 2024). A notable application of LLMs is in question-answering systems, including chatbots, which are prompted with relevant, but unstructured information from knowledge bases to generate accurate answers—an ability we refer to as knowledge comprehension. Knowledge comprehension is a crucial aspect of language understanding, typically assessed in human learners across all education levels (Bloom, 1956; National Center for Education Statistics, 2024; Educational Testing Service, 2024; IDP IELTS, 2024). However, despite being envisioned as superhuman agents (Xi et al., 2023), SOTA LLMs often struggle with basic knowledge comprehension tasks in real-world applications. For example, in the adjacent figure, Gemini-1.5-Flash gets confused with the in-context information and provides an incorrect and potentially fatal response in a critical scenario.
Given the widespread relevance of knowledge comprehension tasks, numerous benchmarking studies have been conducted to evaluate LLMs in this domain (Liang et al., 2023; Chen et al., 2021; Yang et al., 2018; Wang et al., 2023; Tang & Yang, 2024). However, their results have been inconsistent. For example, Wei et al. (2023) demonstrates that larger models are robust to label noise in test sets, while Shi et al. (2024) argues that such models can become distracted by this noise. Conversely, Olsson et al. (2022) indicates that models exhibit similar induction heads for in-context learning, regardless of their size. These discrepancies can be attributed to the following inherent limitations in the nature of these evaluations.
Firstly, LLMs are trained on all available texts, including the standard benchmarking datasets. As a result, models may have encountered the exact prompts used in some test sets, leading to test set leakage during evaluation (Mirzadeh et al., 2024). Consequently, evaluations relying on these test sets become unreliable. Secondly, even in cases where the benchmark studies construct their own datasets, these evaluations remain empirical and assess only a limited subset of potential inputs from the vast space of possible knowledge comprehension prompts. Thirdly, for deployment in practical scenarios, it is also crucial for the evaluations to be comprehensive and reliable - potentially with a guarantee for their validity. Therefore, the following question arises: How can we assess LLM performance in knowledge comprehension tasks, ensuring reliability of accuracy and generalizability?
We argue that to reliably deploy LLMs in safety-critical applications, their empirical evaluation on limited test sets is not sufficient. Instead, they must be evaluated on diverse and large sets of prompts (consisting of different querying styles, large in-context information that can distract models like above, noise, etc.) for knowledge comprehension. Further instead of naively creating prohibitively large test sets, we propose to use precise mathematical representations of these prompts, aka specifications, to enable a computationally tractable evaluation of LLMs. For reliability of the evaluations, it is desirable to have guarantees on their validity. Although prior works (Seshia et al., 2018; Singh et al., 2019; Shi et al., 2020; Bonaert et al., 2021) have proposed formal specifications and certification with guarantees as a method to evaluate traditional neural networks, we cannot directly use them. This is because there exist no formal specifications for knowledge comprehension. Also, certifying LLMs is hard due to a high number of model parameters and specialized nonlinearities, for which traditional certifiers would lose significant precision, leading to inconclusive analysis.
Our approach. Our work proposes a novel formal specification and certification framework, LLMCert-C (LLM Certification of Knowledge Comprehension) that provides reliable performance assessment of the knowledge comprehension capabilities of LLMs. To encode meaningful and diverse knowledge comprehension prompts, we rely on a structured representation of ground truth knowledge - a knowledge graph. This representation enables us to define new probabilistic specifications, where each specification represents a distribution over knowledge comprehension prompts. The prompt distributions, derived from knowledge graphs, are generated with prohibitively large support, consisting of practical prompts with queries over the knowledge graph. These prompts contain substantial amounts of unstructured, relevant information, along with naturally occurring noise—such as distracting texts and information shuffling—that can influence the LLM’s responses. A certificate for a given specification quantifies the probability that the LLM provides a correct response to any prompt sampled from the specification distribution.
LLMCert-C employs a quantitative certification approach that overcomes the limitations of traditional methods for LLM architectures by using a query-based method relying on input-output examples. This also enables using LLMCert-C for closed-source, API-accessible models. In contrast to quantitative certification, binary certificates (Gehr et al., 2018; Wang et al., 2021a) for specification compliance can be trivially invalidated due to the ease of constructing failure examples where the desirable property does not hold (Xu et al., 2024; Vega et al., 2023). LLMCert-C generates probabilistic formal guarantees consisting of high-confidence bounds on the probability of correct knowledge comprehension. LLMCert-C leverages binomial proportion confidence intervals (Clopper & Pearson, 1934) to generate high-confidence, tight certification bounds to estimate the underlying probability measure. The use of bounds is particularly advantageous as they also account for the inherent uncertainty in the estimation. We certify with 2 practical knowledge graphs - PrimeKG (Chandak et al., 2023) over precision medicine knowledge and Wikidata5m (Wang et al., 2021b) over general knowledge. We use the generated certificates to establish novel performance hierarchies with formal guarantees among SOTA LLMs. We also show a consistent decline in LLM performance due to noise in prompts. While formal analysis has been conducted on individual generations of LLMs (Quach et al., 2024) and on counterfactual bias (Chaudhary et al., 2024) in prior work, there is no analysis for the average-case performance of LLMs in knowledge comprehension. Figure 1 gives an overview of LLMCert-C.
Main Contributions:
@@
    We define knowledge comprehension desirable from LLM responses as a formal specification, using knowledge graphs. We develop novel specifications over large sets of prompts with natural noise, such as distracting text and shuffled information order. We generate and experiment with specifications in safety-critical applications in precision medicine and general question-answering using popular knowledge graphs such as PrimeKG and Wikidata5m.
@@
    We model certification of any target LLM with query-access for a given specification as a probability estimation problem and leverage confidence intervals to generate high-confidence guaranteed bounds on the probability of correct knowledge comprehension. We open-source our code at https://github.com/uiuc-focal-lab/LLMCert-C and provide a note to practitioners in Appendix A.
@@
    We observe that as model parameters increase, knowledge comprehension in LLMs improves with high confidence. We show a consistent decline in SOTA LLMs’ performance at answering practical prompts with natural noise.
@@ 2 Certifying knowledge comprehension
Knowledge comprehension is the ability of a model to retrieve and reason over relevant information from context or parametric knowledge to answer questions succinctly and correctly, according to the provided context. Figure 1 gives an overview of our knowledge comprehension certification framework, LLMCert-C. Our framework is agnostic to the internal structure of the target model ${\mathcal{L}}$ allowing any text-to-text generating model.
Figure 1: Overview of LLMCert-C. (a) A KG ${\mathcal{G}}$. (b) A randomly chosen path DAG ${\mathcal{G}}^{\prime}$ originating at fixed node ‘Paracetamol Overdose’, from the various other possible DAGs from node in ${\mathcal{G}}$. (c) A prompt created from ${\mathcal{G}}^{\prime}$ having contexts of the nodes in ${\mathcal{G}}^{\prime}$, a distractor context (highlighted in orange, as the node for ‘Vomiting’ is a distractor for ${\mathcal{G}}^{\prime}$), and a query from ${\mathcal{G}}^{\prime}$. (d) The target LLM’s response to the prompt, validated using the correct answer. (e) Certifier obtains bounds on the probability of correct response using $n$ samples of LLM responses to randomly sampled prompts.
@@ 2.1 Specifying knowledge comprehension
We formally define knowledge comprehension using a knowledge graph (KG, §B.1) ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$. The adjacent grammar defines a KG. Let $\mathcal{V}$ denote the vocabulary of ${\mathcal{L}}$’s tokens. $\mathcal{V}^{+}$ denotes the set of concatenation of non-empty sequences of elements of $\mathcal{V}$.
  KG Grammar   1. $\gamma$    := $\mathcal{V}^{+}$ 2. $\eta$    := $\mathcal{V}^{+}$ 3. $\mathcal{A}$    := $[\eta_{1},\eta_{2},\dots]$ 4. $v$    := $(\gamma,\mathcal{A})$ 5. $e$    := $((v_{1},v_{2}),\mathcal{A})$ 6. $\mathcal{N}$    := $[v_{1},v_{2},\dots]$ 7. $\mathcal{E}$    := $[e_{1},e_{2},\dots]$ 8. $\mathcal{G}$    := $(\mathcal{N},\mathcal{E})$
Let each node $v$ of ${\mathcal{G}}$ (line 2.1) consist of a finite list of synonymous names (a.k.a. aliases, $\mathcal{A}$) that can be used to refer to the node, and a context $\gamma$ that provides more information about the node and its relations with other nodes.
Let $v_{\mathcal{A}}$ and $v_{\gamma}$ denote the aliases and context of $v$ respectively. For example, in Wikidata5m (Wang et al., 2021b), each node has aliases consisting of the identifiers mentioned for the subject of the corresponding Wikipedia page and context which is the abstract of the page. Each (directed) edge $e$ in ${\mathcal{G}}$ (line 2.1) is an ordered pair of related nodes where the relation is identified by a set of synonymous aliases $\mathcal{A}$ for the edge. Let $(v_{1},v_{2})$ denote any edge from node $v_{1}$ to $v_{2}$ in ${\mathcal{G}}$, where $(v_{1},v_{2})_{\mathcal{A}}$ denotes the aliases of the edge. ${\mathcal{G}}$ (line 2.1) is a finite collection of nodes ${\mathcal{N}}$ and edges ${\mathcal{E}}$.
A directed acyclic sub-graph (DAG) ${\mathcal{G}}^{\prime}=({\mathcal{N}}^{\prime},{\mathcal{E}}^{\prime})$ of ${\mathcal{G}}$ is a subgraph of ${\mathcal{G}}$ (${\mathcal{N}}^{\prime}\subseteq{\mathcal{N}},{\mathcal{E}}^{\prime}\subseteq{\mathcal{E}}$) consisting of directed edges and with no cycles. Let the nodes of ${\mathcal{G}}^{\prime}$ be topologically ordered as ${\mathcal{N}}^{\prime}=[v_{1},v_{2}\dots,v_{l}]$. Let the $i^{th}$ ($i\in[1,l]$) nodes of ${\mathcal{N}}^{\prime}$ from $v_{1}$ and backwards from $v_{l}$ be ${\mathcal{N}}^{\prime}[i]\coloneqq v_{i}$ and ${\mathcal{N}}^{\prime}[-i]\coloneqq v_{l-i+1}$ respectively. Nodes ${\mathcal{N}}^{\prime}_{source}$ with no incoming edges are source nodes of ${\mathcal{G}}^{\prime}$. Nodes ${\mathcal{N}}^{\prime}_{sink}$ with no outgoing edges are sink nodes of ${\mathcal{G}}^{\prime}$. DAGs generalize paths that are typically used to define multi-hop knowledge comprehension queries (Yang et al., 2018; Welbl et al., 2018). Hence, we define multi-hop knowledge comprehension queries with DAGs in Definition 2.1.
@@
(Multi-hop queries). Consider a DAG ${\mathcal{G}}^{\prime}$ in ${\mathcal{G}}$ with nodes ${\mathcal{N}}^{\prime}$. A multi-hop query ${\mathcal{Q}}$ for ${\mathcal{N}}^{\prime}$ is to identify the sink nodes ${\mathcal{N}}^{\prime}_{sink}$, given aliases of source nodes ${\mathcal{N}}^{\prime}_{source}$ and aliases of all edges in DAG ${\mathcal{G}}^{\prime}$. To avoid ambiguity regarding correct answer, we enforce ${\mathcal{N}}^{\prime}_{sink}=\{{\mathcal{N}}^{\prime}[-1]\}$ in DAGs for multi-hop knowledge comprehension queries.
The ${\mathcal{G}}$ naturally encodes several multi-hop queries, that form the specification distributions for a language model ${\mathcal{L}}$. To aid reasoning and reduce dependence on parametric knowledge, prompts include alongside relevant textual context needed to identify the intermediate and final nodes. The context may contain natural noise, such as distractor texts or jumbled information, may occur in practical settings. We model such noise in our prompt distributions as described next.
2.1.1 Natural noise in prompts
Distractors. Prior works (Shi et al., 2023) on analyzing reasoning in Language Models (LMs) have shown the negative influence of irrelevant information (distractor) in prompts on LM performance. Hence, we include distractor texts in prompt contexts and specify that model response should not use the distractor information. Contexts of nodes $\tilde{v}$ adjacent to any node ${\mathcal{N}}[i]$ ($i\in[1,l]$) in ${\mathcal{N}}$ that are not the sink node or its predecessors, such that the aliases of the relation $({\mathcal{N}}[i],\tilde{v})$ are the same as those of $({\mathcal{N}}[i],{\mathcal{N}}[j])$, $j>i$ in the DAG, can serve as effective distractors for ${\mathcal{Q}}$ (Definition 2.2). This is because, at any intermediate step, the LM ${\mathcal{L}}$ can pick $\tilde{v}$ as the response, which can deviate ${\mathcal{L}}$’s reasoning from ${\mathcal{N}}$. Nodes adjacent to ${\mathcal{N}}[-1]$ and its predecessors in ${\mathcal{N}}$ are not distractors. For the former, the LM must have already reached the final answer before reaching its adjacent nodes, hence answering ${\mathcal{Q}}$. In the latter, adjacent nodes following same relation are valid answers, not distractors.